# Information over-squashing in language tasks

### v1

@@ 3.1 Copying
Figure 3: Gemini 1.5 being prompted to (i) sum $1+\dots+1$ (left),(ii) Count the number of ones in a sequence of 1s (center), and (iii) Count the number of ones in a sequence of ones and zeroes (the sequence is a Bernoulli sequence with probability of sampling a one being 0.7) (right).
@@ 3.2 Counting
Figure 4: Frequency of different outputted values for Gemini 1.5 for the counting tasks.
Results are presented in Figure 3. It is clear that the performance rapidly deteriorates with the sequence length. It is also interesting to see that the error seems to increase with the sequence very rapidly. For instance in task (i), the LLM is quite likely to predict the value of ‘100’ once the sequence reaches a size around or larger than $100$. Such an observation provides motivating evidence for the argument that Transformers may not be in fact mechanically counting but rather perform a type of crude subitizing. This explains why arguably ‘common’ numbers such as $100$ are much more likely to be outputted by the LLM and why in tasks such as (i) and (ii) the values near $100$ have relatively lower error. This does not happen in task (iii) as the response should actually be around $70\%$ of the sequence length due to the sequence sampling procedure, explaining why the absolute error actually seems to increase around a sequence length of $100$. Figure 4 further showcases this issue, more clearly showing how $100$ is by far the most common response.
@@ 4 Representational Collapse
Let $\mathbf{v}^{(0)}\in\mathbb{R}^{n\times d}$ be a sequence and $\mathbf{v}^{*(0)}\in\mathbb{R}^{(n+1)\times d}$ be another sequence equal to $\mathbf{v}^{(0)}$ with the last token of $\mathbf{v}^{(0)}$ repeated. Assume that the positional encoding information decays to 00 with the distance, e.g. as in RoPE [23]. Then, their representations become arbitrarily close as $n$ increases.
@@ 6 Counting
Counting in certain situations becomes impossible due to representational collapse.
@@ 8 Acknowledgements
We would like to thank Wojciech Marian Czarnecki (Google DeepMind), Simon Osindero (Google DeepMind), and Timothy Nguyen (Google DeepMind) for the valuable comments and suggestions regarding this work. We also thank Constantin Kogler (University of Oxford) for helping to strengthen the theory on representational collapse.
@@ B.1 Representational Collapse
We now study how the total variation between two softmax distributions behaves in the limit of the sequence length. In particular, we show that if the positional encoding information decays to 00 with the sequence length, e.g. as in RoPE [23], then the total variation between the two sequences goes to 00 with $n$. Such a result is presented in Lemma B.2
@@
$\leq\frac{\epsilon\min(Z,Z^{*})}{ZZ^{*}}+\frac{\epsilon\min(Z,Z^{*})}{ZZ^{*}}$
@@
We are now ready to show the main result on representational collapse. In particular, we show that given two sequences of length $n$ and $n+1$ where the second sequence is the same as the first with a final repeated token, their representations become arbitrarily close. Importantly, we require that the information from the positional encodings decays to 00 as the distance grows between tokens, e.g. as in RoPE. The final token repetition is important as otherwise the residual connection in the Transformer would not make the representations necessarily converge.
@@
Let $\mathbf{x}\in\mathbb{R}^{n-1\times d}$ be an underlying growing token sequence. Let $\mathbf{v}^{(0)}=[\mathbf{v}\ \mathbf{v}_{a}]^{T}\in\mathbb{R}^{n\times d}$ and $\mathbf{v}^{*(0)}=[\mathbf{v}\ \mathbf{v}_{a}\ \mathbf{v}_{a}]^{T}\in\mathbb{R}^{n+1\times d}$ be two sequences for a final repeated token $\mathbf{x}_{a}\in\mathbb{R}^{d}$, with all token representations bounded. Further, assume that the positional encodings decay with distance to 00, e.g. as in RoPE. Then, for large enough $n\in\mathbb{N}^{+}$, we have that the representations are under any $\epsilon$:

### v2

@@ 3.1 Copying
Figure 3: Gemini 1.5 being prompted to sum $1+\dots+1$ (Column 1), Count the number of ones in a sequence of 1s (Column 2), Count the number of ones in a sequence of ones and zeroes (the sequence is a Bernoulli sequence with probability of sampling a one being 0.7) (Column 3), and to counter the number of times a word appears in a sentence (Column 4).
@@ 3.2 Counting
Figure 4: Frequency of different outputted values for Gemini 1.5 for the counting tasks. The large density at $100$ suggests that Gemini is likely not counting, but instead possibly performing some crude form of subitising.
Results are presented in Figure 3. It is clear that the performance rapidly deteriorates with the sequence length. It is also interesting to see that the error seems to increase with the sequence very rapidly. For instance in task (i), the LLM is quite likely to predict the value of ‘100’ once the sequence reaches a size around or larger than $100$. Such an observation provides motivating evidence for the argument that Transformers may not be in fact mechanically counting but rather perform a type of crude subitising. This explains why arguably ‘common’ numbers such as $100$ are much more likely to be outputted by the LLM and why in tasks such as (i) and (ii) the values near $100$ have relatively lower error. This does not happen in task (iii) as the response should actually be around $70\%$ of the sequence length due to the sequence sampling procedure, explaining why the absolute error actually seems to increase around a sequence length of $100$. Figure 4 further showcases this issue, more clearly showing how $100$ is by far the most common response.
@@ 4 Representational Collapse
Let $\mathbf{v}^{(0)}\in\mathbb{R}^{n\times d}$ be a sequence and $\mathbf{v}^{*(0)}\in\mathbb{R}^{(n+1)\times d}$ be another sequence equal to $\mathbf{v}^{(0)}$ with the last token of $\mathbf{v}^{(0)}$ repeated. Assume that the positional encoding information decays to 00 with the distance. Then, their representations become arbitrarily close as $n$ increases.
@@ 6 Counting
Counting in certain situations becomes impossible due to representational collapse and finite floating point precision.
Corollary 6.2 shows how our main result on representational collapse points to practical issues when it comes to certain styles of prompts. When paired with low floating point arithmetic precision, representation collapse becomes problematic.
@@ 8 Acknowledgements
We would like to thank Wojciech Marian Czarnecki (Google DeepMind), Simon Osindero (Google DeepMind), and Timothy Nguyen (Google DeepMind) for the valuable comments and suggestions regarding this work. We also thank Constantin Kogler (University of Oxford) for providing useful insights on the theoretical aspects of the work.
@@ B.1 Representational Collapse
We now study how the total variation between two softmax distributions behaves in the limit of the sequence length. In particular, we show that if the positional encoding information decays to 00 with the sequence length, then the total variation between the two sequences goes to 00 with $n$. Such a result is presented in Lemma B.2
@@
$=\sum_{i=1}^{n}\frac{\left\lvert Z^{*}-Z\right\rvert e^{\mathbf{x}_{i}}}{ZZ^{*}}+\frac{Z\lvert e^{\mathbf{x}_{i}}-e^{\mathbf{x}^{*}_{i}}\rvert}{ZZ^{*}}$
$=\frac{\left\lvert Z^{*}-Z\right\rvert}{ZZ^{*}}Z+\sum_{i=1}^{n}\frac{\lvert e^{\mathbf{x}_{i}}-e^{\mathbf{x}^{*}_{i}}\rvert}{Z^{*}}$
$\leq\frac{\epsilon\min(Z,Z^{*})}{Z^{*}}+\frac{\epsilon\min(Z,Z^{*})}{Z^{*}}$
@@
We are now ready to show the main result on representational collapse. In particular, we show that given two sequences of length $n$ and $n+1$ where the second sequence is the same as the first with a final repeated token, their representations become arbitrarily close. Importantly, we require that the information from the positional encodings decays to 00 as the distance grows between tokens. The final token repetition is important as otherwise the residual connection in the Transformer would not make the representations necessarily converge.
@@
Let $\mathbf{x}\in\mathbb{R}^{n-1\times d}$ be an underlying growing token sequence. Let $\mathbf{v}^{(0)}=[\mathbf{v}\ \mathbf{v}_{a}]^{T}\in\mathbb{R}^{n\times d}$ and $\mathbf{v}^{*(0)}=[\mathbf{v}\ \mathbf{v}_{a}\ \mathbf{v}_{a}]^{T}\in\mathbb{R}^{n+1\times d}$ be two sequences for a final repeated token $\mathbf{x}_{a}\in\mathbb{R}^{d}$, with all token representations bounded. Further, assume that the positional encodings decay with distance to 00. Then, for large enough $n\in\mathbb{N}^{+}$, we have that the representations are under any $\epsilon$:
@@ C.3 Synthetic Experiments on Representational Collapse
C.4 Effect of Positional Encodings
We include a synthetic experiment where we verify the occurrence of the representational collapse phenomenon with different Positional Encodings, namely: Alibi [24], the original Absolute Positional Encodings (APE) [30], and No Positional Encodings (NoPE) [15]. In our synthetic experiment, we sample $n$ queries and keys independently directly from a standard Gaussian. We then construct a related sequence of length $n+1$ by repeating its last element. We report the $L_{1}$ distance between the two sequences after a single decoder-Transformer layer, as done for the other representational collapse experiments. We consider a Transformer with a hidden dimension of $64$, a single attention head, and we apply normalisations to simulate layer norm. The Transformer is not trained, but only used to simulate the propagation of information of queries and keys sampled from a Gaussian distribution. The results are shown in Figure 10. Representational collapse seems to occur with all $4$ positional encodings, with the convergence of the representations happening at a similar sequence lengths.
We would like to highlight that our condition on the decay of RoPE necessary to fulfill the requirement of Theorem 4.2 is inspired by claims of the decay of RoPE coming from the original work by Su et al. [26]. However, recent work has shown that such claims may not be strictly always upheld [4], as the original claims relied on very specific conditions on the queries and keys. We are of the opinion; however, that the range of synthetic and real-world experiments in this work support our representational collapse claims in practice with RoPE. A more precise mathematical treatment of RoPE specifically is therefore left as future work.
Figure 10: We sample $n$ queries, keys, and values independently from a standard Gaussian, applying different positional encodings. We then construct sequences of length $n+1$, by repeating the $n$-th token. We report the $L_{1}$ distance between the last tokens of the sequences of length $n$ and $n+1$ after one decoder-only Transformer layer. We set the hidden dimension to $64$, use a single attention head, and normalise appropriately to simulate the effects of LayerNorm. The y-axis is shown in log-scale.
C.5 Ablation on Prompt Structure
We ablate the prompt structure specifically for the copying task. In particular, we consider the prompts: (Type 1) “What is the last digit of the following sequence? Please answer exactly as ‘The answer to your question is: <ANSWER>’ ”. Here is the sequence: {seq} and (Type 2) “Please answer exactly as ‘The answer to your question is: <ANSWER>’. What is the last digit of the following sequence? {seq}”. The results are presented in Figure 11. We find that the prompt indeed does affect the performance on the task, as the prompt affects the distribution of the attention over the layer. However, for both types of prompts, the model ends up failing, in accordance with our theory. We also show for completeness, in Figure 12, that representational collapse occurs in Gemma 7B also for the ‘Type 1’ prompt.
Figure 11: Performance of a Gemini model on the following prompts: (Type 1) “What is the last digit of the following sequence? Please answer exactly as ‘The answer to your question is: <ANSWER>’ ”. Here is the sequence: {seq} and (Type 2) “Please answer exactly as ‘The answer to your question is: <ANSWER>’. What is the last digit of the following sequence? {seq}”
Figure 12: Representational collapse in Gemma for the prompt: “What is the last digit of the following sequence? Please answer exactly as ‘The answer to your question is: <ANSWER>’ ”. Here is the sequence: {seq} and (Type 2) ”
C.6 Local sliding window attention
A fundamental limitation of an attention mechanism that leverages the softmax function is that it cannot remain sharp, especially as the sequence length grows [31]. This is in fact a key intuition that we exploit to show our result on representational collapse. A good way to address representational collapse and the related phenomenon of over-squashing is then that of limiting the spread of the softmax function, by directly limiting the amount of tokens the attention mechanism pays attention to. This mechanism is often referred to as a local sliding window and is a major architectural change present in Gemma 2 [28]. We believe that such an architectural change elegantly addresses representational collapse and over-squashing at the source as it avoids the issues that come with growing token sequences – something which our theory often exploits.